OpenStudy (anonymous):

solve for t: 17e^(0.07t)=11e^(0.05t) I have so far gotten to this step: 17t ln(0.07)=tln(0.05)

OpenStudy (anonymous):

HELP ME PLEASE

OpenStudy (anonymous):

divide by e^(0.05t) on both sides

OpenStudy (anonymous):

and post what you get.

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

17tln(0.07)-tln(0.05)=11

OpenStudy (anonymous):

\[e ^{x} / e ^{y} = e ^{x-y}\]\[\ln{e^{x-y}} = x - y\]

OpenStudy (anonymous):

how did you get to the step that you got to? what did you do to get 17t ln(0.07)=tln(0.05)? because that is incorrect.

OpenStudy (anonymous):

I did this : 17e^(0.07t)/(e^(0.05t) = 11

OpenStudy (anonymous):

ok, thar part is right. what did you do after that?

OpenStudy (anonymous):

okay, and then?

OpenStudy (anonymous):

is it possible that you can take a picture of your notes and attach it here? that way it is much faster.

OpenStudy (anonymous):

according to the quotient rule u subtract?

OpenStudy (anonymous):

right, you subtract

OpenStudy (anonymous):

so what did you get?

OpenStudy (anonymous):

17tln(0.07) - tln(0.05)=11

OpenStudy (anonymous):

why did you do tln0.07-tlon0.05? what rule is that?

OpenStudy (anonymous):

idk. can u show me the next step. plz

OpenStudy (anonymous):

alright so 17e^0.07t = 11e^0.05t Step 1: divide both sides by e^0.05t ...so you have (17e^0.07t)/(e^0.05t) = 11 Step 2: since e^x/y = e^x-y ...you can make this 17e^(0.07t-0.05t) =11 ...which equals???? 17e^0.02t = 11 Step 3: divide 17 out of there ...so e^0.02t = (11/17) step 4: take the natural log of both sides ...so ln e^0.02t = ln (11/17) step 5: e goes away now and you have 0.02t = ln (11/17) step 6: divide both sides by 0.02 so now it is t = (ln (11/17)/(0.02)) SO NOW..... t = -21.76590356 Check to see if this works in the original equation

OpenStudy (anonymous):

hold on let me work it out

OpenStudy (anonymous):

yup it works. i understand now, i was not certain at first if u can use the e^x/y rule

OpenStudy (anonymous):

i understand the simplification

OpenStudy (anonymous):

i messed up in my steps by going too fast and said e^x/y = e^x-y i meant to say (e^x) / (e^y) = e^x-y but i think you already figured that out

OpenStudy (anonymous):

the rule is e^x/e^y = e^(x-y), not e^x/y = e^(x-y)

OpenStudy (anonymous):

yep...caught myself there :)

OpenStudy (anonymous):

yup. qoptient rules for natural logs right?

OpenStudy (anonymous):

lol. no its the quotient rule for ALL fractions.

OpenStudy (anonymous):

yep....now if you'll excuse me, i have to go make some natural logs of my own....;o

OpenStudy (anonymous):

TMI

OpenStudy (anonymous):

hahaha

OpenStudy (anonymous):

suprisingly im an engineering major

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

surprisingly, i earned a bachelor's in biochemistry....when i wasnt drunk...